# Observation, Analysis, and Solution: Exploring Strong Lightweight Vision Transformers via Masked Image Modeling Pre-Training

### v1

@@ 3.2 Fine-Tuning Evaluation on ImageNet Classification
Flowers Pets Aircraft Cars CIFAR100 iNat18 ADE20K LaSOT COCO (det./seg.)
@@
Observation 2. The above observation encourages us to further explore how the enhanced lightweight ViTs perform in comparison to the recent SOTA lightweight ViT derivatives (including both the hybrid and non-hybrid) as introduced in Tab. VII. It can be observed that the enhanced vanilla ViT-Tiny through MIM pre-training can also achieve SOTA performance on par with some superior ViT derivatives with comparable parameter number. It is noteworthy that these delicately designed architectures in the lightweight regime, despite lack of pre-training, achieve top performance on the ImageNet classification task by seeking to (1) extend the supervised fine-tuning duration to a much longer schedule (e.g., 1000 epochs in LeViT [17] and PiT [19]), (2) introduce distillation-driven training to regularize the fine-tuning (e.g., in LeViT [17], XCiT [18], PiT [19], and CaiT [20], etc.), or (3) use some recent techniques like relative position bias [116, 117, 118, 119, 120] (e.g., in LeViT [17] and Swin [86]), conditional positional encoding [122] (e.g., in EdgeViT [23]) and LayerScale (e.g., in XCiT [18]). The competing results of the enhanced vanilla ViT-Tiny demonstrate the benefits of studying MIM pre-training on extremely simple lightweight ViTs, which may be orthogonal to the network architecture design methodology in the ViT derivatives. Besides, extremely simple architectures are usually friendly to the real-world deployment by getting rid of the complicated modules existing in some ViT derivatives (e.g., EdgeViT-XS [23], MobileViT-S [21], Mobile-Former-294M [22], SeaFormer-L [27]).
Observation 3. We also empirically examine whether the self-supervised pre-training on lightweight ViTs can benefit from large-scale pre-training data. In Tab. II, we study the effect of varying the pre-training data size and class distribution with two typical SSL methods (i.e., MoCo-v3 and MAE) applied on ViT-Tiny. We first consider the much larger dataset IN21K, however, few improvements are observed for both MoCo-v3 and MAE. We believe the limited capacity of extremely simple lightweight ViTs also bottlenecks the representation quality of self-supervised pre-training even with the “LLM-like” scaling of data. We further consider two subsets of IN1K containing 10% and 1% of the total examples (IN1K 10% and IN1K 1%) balanced in terms of classes [123] and one subset with long-tailed class distribution [124] (IN1K-LT). Surprisingly, MAE pre-training has marginal performance declines on these subsets, showing more robustness than MoCo-v3 in terms of the pre-training data size and class distribution.
3.3 Transfer Learning Evaluation on Downstream Tasks
We further examine the transfer learning performance of the pre-trained models in Tab. III. We only consider the models pre-trained based on IN1K for simplicity, involving their transfer performance on some other downstream classification tasks and dense prediction tasks.
For the classification tasks, we investigate the performance on the relatively smaller datasets than IN1K to question if the pre-trained models transfer well on these data-insufficient tasks, including Flowers-102 [125] (Flowers for short), Oxford-IIIT Pets [126] (Pets), FGVC-Aircraft [127] (Aircraft), Stanford Cars [128] (Cars), CIFAR100 [129], iNaturalist 2018 [130] (iNat18). For all these datasets except iNat18, we conduct a small sweep for some key fine-tuning parameters, i.e., the learning rate, training epoch number and layer-wise lr decay coefficient, to achieve the optimal fine-tuning results on these datasets (see Sec. A.2 for more details). Besides, we adopt random resized crop and random horizontal flipping as the augmentations and do not use any regularization (e.g., weight decay, dropout, or the stochastic depth regularization technique [131]). For iNat18, we follow the same training configurations to those on ImageNet (see Sec. A.1).
@@ 3.3 Transfer Learning Evaluation on Downstream Tasks
Figure 3: Layer representation similarity as heatmaps for the investigated ViT-Tiny models from different pre-training methods, with x and y axes indexing the layers (the 0 index indicates the patch embedding layer), and higher values indicate higher similarity. The fully-supervised baseline based on our recipe (see Tab. I) on IN1K is used as the reference.
Observation 4. As shown in Tab. III, we first find it hard to achieve promising transfer evaluation results for our modified ViT-Tiny on the evaluated downstream tasks by training from scratch without any pre-training. The results significantly lag behind those models that adopt various pre-training. Second, despite that using various pre-training methods shows better performance, the relative superiority and inferiority comparisons between these pre-training methods exhibit distinct characteristics from those on ImageNet (see Tab. I). We find that downstream data scale matters during transfer learning. The MIM-based pre-training approaches tend to achieve downstream performance inferior to the fully-supervised counterpart, while the performance gap is narrowed more or less as the data scale of the downstream task increases. Moreover, MIM-based pre-training even shows inferior results to CL-based approaches. We conjecture that it is due to their different layer behaviors during pre-training and fine-tuning as discussed in detail in the following section.
4 Analysis
In this section, we introduce linear probing evaluation and some model analysis methods to reveal the secrets of pre-training, study the pattern of layer behaviors during pre-training and fine-tuning, and investigate what matters for downstream performances. For the sake of simplicity, we only investigate the ImageNet classification task and other downstream classification tasks shown in Tab. III and use the supervised pre-training baseline, the CL-based entries DINO [1] and MoCo-v3 [6], the best MIM-based entry MaskFeat [32], and the worst two MIM-based entries MAE [12] and SimMIM [28] in Tab. III for analysis in this section.
4.1 Linear Probing Evaluation
We first learn linear classifiers on frozen features for different classification tasks in our linear probing evaluation. It reflects how the representations obtained by the pre-trained models are linearly separable w.r.t. semantic categories. As shown in Tab. IV, the linear probing performance is consistently lower than the fine-tuning performance. Furthermore, while linear probing evaluation is informative for exploiting the pre-trained frozen models, it may not correspond to which models perform best after fine-tuned, especially for those downstream tasks with relatively sufficient labeled data, e.g., iNat18 and IN1K. In other words, it may lead to an underestimation of the value of some pre-trained models in the practical utility [12, 121] on downstream tasks. We attribute it to that linear probing only evaluates the final representation of the pre-trained models, which makes it overlook the value of providing good initialization for lower layers.
TABLE IV: Linear Probing Evaluation of Pre-Trained Models on the Classification Tasks. Top-1 Accuracy Is Reported.
Methods Datasets
Flowers Pets Aircraft Cars CIFAR100 iNat18 IN1K
Supervised
Our Recipe 91.0 92.0 41.2 47.9 73.6 39.8 -
CL
DINO 95.0 90.9 53.5 53.6 74.9 44.0 66.1
MoCo-v3 93.2 83.5 44.8 44.5 73.4 36.2 62.1
MIM
MaskFeat 65.5 34.2 14.9 13.8 39.6 3.9 23.1
MAE 48.9 25.0 12.8 8.8 31.0 1.4 23.3
SimMIM 44.1 18.5 11.2 6.6 23.3 0.9 14.8
@@ 4.1 Linear Probing Evaluation
From another point of view, we think the failure of MIM-based self-supervised pre-training to perform better than CL-based models for the data-insufficient downstream tasks (as shown in Tab. III) is due to the reason that CL pre-training learns more semantics at an abstract level relevant to recognition in higher layers than MIM pre-training, which also helps the frozen models from CL pre-training to achieve better linear probing evaluation results than MIM (as shown in Tab. IV). This analysis is already evident in the CL-based DINO [1] as its pre-training results show that the class-specific features can be automatically learned, even leading to unsupervised object segmentation. This motivates us to further analyze the layer-wise representations and attention maps across and within different models to delve into the working mechanism of the different SSL pre-training methods for lightweight ViTs.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
4.2 Layer-Wise Representation Analysis
We adopt the Centered Kernel Alignment (CKA) metric [^3] https://github.com/AntixK/PyTorch-Model-Compare [138, 139] to analyze the layer-wise representation similarity across and within different models. Specifically, CKA takes two feature maps (or representations) $\bm{X}$ and $\bm{Y}$ as input and computes their normalized similarity $S(\bm{X},\bm{Y})$ in terms of the Hilbert-Schmidt Independence Criterion (HSIC [140]) as
$S(\bm{X},\bm{Y})=\mathrm{CKA}(\bm{K},\bm{L})=\frac{\mathrm{HSIC}(\bm{K},\bm{L})}{\sqrt{\mathrm{HSIC}(\bm{K},\bm{K})\mathrm{HSIC}(\bm{L},\bm{L})}},$ (1)
Analysis 1. We use heatmaps to visualize the layer-wise representation similarity of investigated pre-trained models with respect to a fully-supervised reference model in Fig. 3. We choose the fully-supervised baseline trained on the IN1K classification task based on our recipe (see Tab. I) as the reference because we consider its representations are more recognition aligned and the higher similarity of the examined model’s layers to this supervised baseline can indicate more relevance to recognition for the examined model. Although the similarity does not directly indicate whether the downstream performance is good, it indeed reflects the pattern of layer representations to a certain extent. The similarity within the supervised baseline is also presented (see top-left of Fig. 3). In Sec. B.1, we also replace the above reference model with some much stronger ViT-Base reference models to test the robustness of our analysis.
We notice that lower layers matter more than higher ones if sufficient downstream data is provided. Specifically, we first observe a relatively high similarity between the models from MIM pre-training and the reference model for lower layers, while much lower similarity for higher layers. It indicates that fewer semantics are extracted from MIM pre-training at a more abstract level in higher layers. In contrast, the models from CL pre-training aligns the reference model well across almost all layers. However, the fine-tuning evaluation on IN1K (see Tab. I) shows that adopting MIM pre-training as the initialization improves the performance more significantly than CL pre-training. Thus, we hypothesize that lower layers matter much more than higher ones for the self-supervised pre-training methods to perform well on the data-sufficient IN1K. In order to verify the hypothesis, we design another experiment by sweeping over the number of reserved consecutive leading blocks of pre-trained models for initialization (the remaining blocks after the reserved blocks are randomly initialized), and then fine-tuning the initialized models on IN1K (for the sake of simplicity, we adopt 100-epoch fine-tuning). The left figure of Fig. 4 shows that reserving only a certain number of leading blocks at the lower end can already achieve a significant performance gain over randomly initializing all the blocks (i.e., totally training from scratch) for both MIM-based MAE and CL-based MoCo-v3 pre-training. Whereas, further reserving higher layers leads to marginal gains for both of them, which demonstrates our hypothesis.
@@ 4.3 Attention Map Analysis
Refer to caption    Refer to caption    Refer to caption    Refer to caption
Refer to caption    Refer to caption    Refer to caption    Refer to caption
Refer to caption    Refer to caption    Refer to caption    Refer to caption
@@ 5.1 Improve MAE Pre-Training Based on Distillation
We herein introduce our distillation process in the MAE pre-training framework (as illustrated in Fig. 9) to showcase the effectiveness of our observation-analysis-solution flow in improving the MIM pre-traning of ViT-Tiny on higher layers. Based on the masked autoencoder framework, we introduce our examined MAE pre-trained ViT-Base model in Fig. 7 as the teacher. During the MAE pre-training of ViT-Tiny, the teacher processes the same visible image patches as the student encoder ViT-Tiny, and two kinds of layer-wise attention-based and representation-based distillation losses can be constructed between the corresponding teacher’s and student’s layers for our investigation. The parameters of the student are updated based on the joint backward gradients from the distillation loss and the original MAE’s reconstruction loss, while the teacher’s parameters remain frozen throughout the pre-training process. The original MAE’s reconstruction loss is constructed by using the output of the last layer of the student encoder as the decoder’s input.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
Distill on attention maps or output representations? We conduct a fine-tuning performance comparison (100 epochs for the sake of simplicity) on IN1K for the utilization of the two kinds of different distillation targets during our MAE pre-training (400 epochs). We test the distillation on the 3rd, 6th, 9th and 12th layers individually, i.e., the distillation losses are constructed between the pairs of layers at $1/4$, $2/4$, $3/4$ and $/4$ 4 depth of the teacher and student respectively. As shown in Fig. 9, distillation on the attention maps of the last transformer blocks promotes the performance most, not only surpassing the same kind of distillation on lower layers, but also contributing to more performance gain in comparison to the distillation on the output representations. It is consistent with the analyses in Sec. 4. On the one hand, the lower layers learn good representations themselves during the MAE pre-training without distillation, and thus distillation on these layers contributes to marginal improvements, though distillation on the output representations directly performs a little better than on the attention maps of these layers. On the other hand, the higher layers rely on a good teacher to guide them to capture rich semantic features, and the attention weights on the higher layers of the teacher exhibit some patterns of behavior with much richer semantic knowledge, which may be the reason for the attention map’s superiority over the representation for distillation on higher layers. Fig. 7 also shows that the representations deteriorate on the 12th layer for the MAE pre-trained ViT-Base.
TABLE V: Sweep Over the Intermediate Layer Index to Which the Decoder for Reconstruction Is Connected. Top-1 Accuracy for Classification Tasks and AP for Detection Tasks Are Reported.
@@
We denote the above best MAE pre-training based on attention map distillation on the last layer as D-MAE. As show in the top-left subfigure of Fig. 10, the CKA analysis for the D-MAE pre-trained ViT-Tiny indicates that the rich semantic knowledge of the teacher has been successfully transferred to the pre-trained lightweight student.
TABLE VI: Distillation Significantly Improves Performance on Various Tasks. Top-1 Accuracy Is Reported for Classification Tasks, mIoU Is for ADE20K Segmentation, AUC Is for LaSOT Tracking, and AP Is for COCO Detection (Det.) and Segmentation (Seg.).
@@
Flowers Pets Aircraft Cars CIFAR100 iNat18 IN1K ADE20K LaSOT COCO (det./seg.)
@@
$\Delta$ to Supervised Pre-train -0.4 -2.7 +6.5 +2.6 -0.6 +1.5 +2.2 +1.8 +2.0 +2.1/+2.0
5.2 Decoupled Distillation Further Improves Pre-Training
Decoupled distillation strategy. As shown in  Fig. 3, the deterioration of representation learning on higher layers in the original MAE pre-training of lightweight ViTs indicates that there is some homogeneity among the representation learning of higher layers during the MAE reconstruction of the RGB pixel targets. That means it is not necessary to rely on the last layer for reconstructing the targets. We thus decouple the distillation task from the MAE reconstruction task by connecting the decoder for reconstruction to other intermediate layers of the student ViT-Tiny instead of the last layer during pre-training in Fig. 9. We conduct a sweep over the intermediate layer index ranging from 2 to 12 for the optimal decoupled distillation. The fine-tuning results on some selected datasets (see Tab. V) show that connecting to the 6th, 8th, or 10th layer can all further improve the performance over the model based on D-MAE (i.e., the column with the encoder layer index of 12 for reconstruction in Tab. V). We use the 8th encoder layer for MAE reconstruction in the subsequent experimental analysis and validation by default, and denote this pre-training with decoupled distillation as D2-MAE. It can be seen in Fig. 10 that the attention weights on the higher layers after D2-MAE pre-training exhibit very similar behavior patterns to the ViT-Base teacher (see Fig. 7).
In Tab. V, we also test the scenario of only relying on the attention map distillation to achieve good representation for the lightweight ViT model, i.e., the MAE reconstruction loss in Fig. 9 is abandoned. It is shown that the experimental results without reconstruction are significantly lower than using both the distillation and reconstruction. However, only relying on the attention map distillation has already slightly surpassed the MAE pre-training without distillation for ViT-Tiny. This further demonstrate the effectiveness and superiority of our observation-analysis-solution flow for developing the attention map distillation-based solution.

### v2

@@ 3.3 Transfer Learning Evaluation on Downstream Tasks
Flowers Pets Aircraft Cars CIFAR100 iNat18 ADE20K LaSOT COCO
@@
For the classification tasks, we investigate the performance on the relatively smaller datasets than IN1K to question if the pre-trained models transfer well on these data-insufficient tasks, including Flowers-102 [129] (Flowers for short), Oxford-IIIT Pets [130] (Pets), FGVC-Aircraft [131] (Aircraft), Stanford Cars [132] (Cars), CIFAR100 [133], iNaturalist 2018 [134] (iNat18). For all these datasets except iNat18, we conduct a small sweep for some key fine-tuning parameters, i.e., the learning rate, training epoch number and layer-wise lr decay coefficient, to achieve the optimal fine-tuning results on these datasets (see Appendix A.2 for more details). Besides, we adopt random resized crop and random horizontal flipping as the augmentations and do not use any regularization (e.g., weight decay, dropout, or the stochastic depth regularization technique [135]). For iNat18, we follow the same training configurations to those on ImageNet (see Appendix A.1).
@@
Observation 4. As shown in Tab. 3, we first find it hard to achieve promising transfer evaluation results for our modified ViT-Tiny on the evaluated downstream tasks by training from scratch without any pre-training. The results significantly lag behind those models that adopt various pre-training. Second, despite that using various pre-training methods shows better performance, the relative superiority and inferiority comparisons between these pre-training methods exhibit distinct characteristics from those on ImageNet (see Tab. 1). We find that downstream data scale matters during transfer learning. The MIM-based pre-training approaches tend to achieve downstream performance inferior to the fully-supervised counterpart, while the performance gap is narrowed more or less as the data scale of the downstream task increases. Moreover, MIM-based pre-training even shows inferior results to CL-based approaches. We conjecture that it is due to their different layer behaviors during pre-training and fine-tuning as discussed in detail in the following section.
4 Analysis
In this section, we introduce linear probing evaluation and some model analysis methods to reveal the secrets of pre-training, study the pattern of layer behaviors during pre-training and fine-tuning, and investigate what matters for downstream performances. For the sake of simplicity, we only investigate the ImageNet classification task and other downstream classification tasks shown in Tab. 3 and use the supervised pre-training baseline, the CL-based entries DINO [1] and MoCo-v3 [6], the best MIM-based entry MaskFeat [32], and the worst two MIM-based entries MAE [12] and SimMIM [28] in Tab. 3 for analysis in this section.
Table 4: Linear Probing Evaluation of Pre-Trained Models on the Classification Tasks. Top-1 Accuracy Is Reported.
Methods Datasets
Flowers Pets Cars CIFAR iNat18 IN1K
Supervised
Our Recipe 91.0 92.0 47.9 73.6 39.8 -
CL
DINO 95.0 90.9 53.6 74.9 44.0 66.1
MoCo-v3 93.2 83.5 44.5 73.4 36.2 62.1
MIM
MaskFeat 65.5 34.2 13.8 39.6 3.9 23.1
MAE 48.9 25.0 8.8 31.0 1.4 23.3
SimMIM 44.1 18.5 6.6 23.3 0.9 14.8
4.1 Linear Probing Evaluation
We first learn linear classifiers on frozen features for different classification tasks in our linear probing evaluation. It reflects how the representations obtained by the pre-trained models are linearly separable w.r.t. semantic categories. As shown in Tab. 4, the linear probing performance is consistently lower than the fine-tuning performance except for DINO on Pets. Furthermore, while linear probing evaluation is informative for exploiting the pre-trained frozen models, it may not correspond to which models perform best after fine-tuned, especially for those downstream tasks with relatively sufficient labeled data, e.g., iNat18 and IN1K. In other words, it may lead to an underestimation of the value of some pre-trained models in the practical utility [12, 125] on downstream tasks. We attribute it to that linear probing only evaluates the final representation of the pre-trained models, making it overlook the value of providing good initialization for lower layers.
@@ 4.1 Linear Probing Evaluation
Figure 3: Layer representation similarity as heatmaps for the investigated ViT-Tiny models from different pre-training methods, with x and y axes indexing the layers (the 0 index indicates the patch embedding layer), and higher values indicate higher similarity. The fully-supervised baseline based on our recipe (see Tab. 1) on IN1K is used as the reference.
@@
From another point of view, we think the failure of MIM-based self-supervised pre-training to perform better than CL-based models for the data-insufficient downstream tasks (as shown in Tab. 3) is due to the reason that CL pre-training learns more semantics at an abstract level relevant to recognition in higher layers than MIM pre-training, which also helps the frozen models from CL pre-training to achieve better linear probing evaluation results than MIM (as shown in Tab. 4). This analysis is already evident in the CL-based DINO [1] as its pre-training results show that the class-specific features can be automatically learned, even leading to unsupervised object segmentation. This motivates us to further analyze the layer-wise representations and attention maps across and within different models to delve into the working mechanism of the different SSL pre-training methods for lightweight ViTs.
4.2 Layer-Wise Representation Analysis
We adopt the Centered Kernel Alignment (CKA) metric [138, 139] to analyze the layer-wise representation similarity across and within different models. Specifically, CKA takes two feature maps (or representations) $\bm{X}$ and $\bm{Y}$ as input and computes their normalized similarity $S(\bm{X},\bm{Y})$ in terms of the Hilbert-Schmidt Independence Criterion (HSIC [140]) as
$S(\bm{X},\bm{Y})$ $=\mathrm{CKA}(\bm{K},\bm{L})$
$=\frac{\mathrm{HSIC}(\bm{K},\bm{L})}{\sqrt{\mathrm{HSIC}(\bm{K},\bm{K})\mathrm{HSIC}(\bm{L},\bm{L})}},$ (1)
Analysis 1. We use heatmaps to visualize the layer-wise representation similarity of investigated pre-trained models with respect to a fully-supervised reference model in Fig. 3. We choose the fully-supervised baseline trained on the IN1K classification task based on our recipe (see Tab. 1) as the reference because we consider its representations are more recognition aligned and the higher similarity of the examined model’s layers to this supervised baseline can indicate more relevance to recognition for the examined model. Although the similarity does not directly indicate whether the downstream performance is good, it indeed reflects the pattern of layer representations to a certain extent. The similarity within the supervised baseline is also presented in Fig. 3. In Appendix B.1, we also replace the reference model with some much stronger ViT-Base models to test the robustness of our analysis.
We notice that lower layers matter more than higher ones if sufficient downstream data is provided. Specifically, we first observe a relatively high similarity between the models from MIM pre-training and the reference model for lower layers, while much lower similarity for higher layers. It indicates that fewer semantics are extracted from MIM pre-training at a more abstract level in higher layers. In contrast, the models from CL pre-training aligns the reference model well across almost all layers. However, the fine-tuning evaluation on IN1K (see Tab. 1) shows that adopting MIM pre-training as the initialization improves the performance more significantly than CL pre-training. Thus, we hypothesize that lower layers matter much more than higher ones for the self-supervised pre-training methods to perform well on the data-sufficient IN1K. In order to verify the hypothesis, we design another experiment by sweeping over the number of reserved consecutive leading blocks of pre-trained models for initialization (the remaining blocks after the reserved blocks are randomly initialized), and then fine-tuning the initialized models on IN1K (for the sake of simplicity, we adopt 100-epoch fine-tuning). The left figure of Fig. 4 shows that reserving only a certain number of leading blocks at the lower end can already achieve a significant performance gain over randomly initializing all the blocks (i.e., training from scratch) for both MAE and MoCo-v3 pre-training. Whereas, further reserving higher layers leads to marginal gains, which demonstrates our hypothesis.
@@ 4.3 Attention Map Analysis
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
5 Solution
@@ 5.1 Improve MAE Pre-Training Based on Distillation
We herein introduce our distillation process in the MAE pre-training framework (as illustrated in Fig. 9) to showcase the effectiveness of our observation-analysis-solution flow in improving the MIM pre-traning of ViT-Tiny on higher layers. We introduce our examined MAE pre-trained ViT-Base model in Fig. 7 as the teacher. During the MAE pre-training of ViT-Tiny, the teacher processes the same visible image patches as the student encoder ViT-Tiny, and two kinds of layer-wise attention-based and representation-based distillation losses can be constructed between the corresponding teacher’s and student’s layers for our investigation. The parameters of the student are updated based on the joint backward gradients from the distillation loss and the original MAE’s reconstruction loss, while the teacher’s parameters remain frozen throughout the pre-training process. The original MAE’s reconstruction loss is constructed by using the output of the last layer of the student encoder as the decoder’s input.
@@
Distill on attention maps or output representations? We conduct a fine-tuning performance comparison (100 epochs for the sake of simplicity) on IN1K for the utilization of the two kinds of different distillation targets during our MAE pre-training (400 epochs). We test the distillation on the 3rd, 6th, 9th and 12th layers individually, i.e., the distillation losses are constructed between the pairs of layers at $1/4$, $2/4$, $3/4$ and $/4$ 4 depth of the teacher and student respectively. As shown in Fig. 9, distillation on the attention maps of the last transformer blocks promotes the performance most, not only surpassing the same kind of distillation on lower layers, but also contributing to more performance gain in comparison to the distillation on the output representations. It is consistent with the analyses in Sec. 4. On the one hand, the lower layers learn good representations themselves during the MAE pre-training without distillation, and thus distillation on these layers contributes to marginal improvements, though distillation on the output representations directly performs a little better than on the attention maps of these layers. On the other hand, the higher layers rely on a good teacher to guide them to capture rich semantic features, and the attention weights on the higher layers of the teacher exhibit some patterns of behavior with much richer semantic knowledge, which may be the reason for the attention map’s superiority over the representation for distillation on higher layers. Fig. 7 also shows that the representations deteriorate on the 12th layer for the MAE pre-trained ViT-Base. We denote the above best MAE pre-training based on attention map distillation on the last layer as D-MAE. As shown in the top-left of Fig. 10, the CKA analysis for the D-MAE pre-trained ViT-Tiny indicates that the rich semantic knowledge of the teacher has been successfully transferred to the lightweight student.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
5.2 Decoupled Distillation Further Improves Pre-Training
Table 5: Sweep over the intermediate layer index to which the decoder for reconstruction is connected. Top-1 accuracy for classification tasks and AP for detection tasks are reported.
@@ 5.2 Decoupled Distillation Further Improves Pre-Training
Decoupled distillation strategy. As shown in  Fig. 3, the deterioration of representation learning on higher layers in the original MAE pre-training of lightweight ViTs indicates that there is some homogeneity among the representation learning of higher layers during the MAE reconstruction of the RGB pixel targets. That means it is not necessary to rely on the last layer for reconstructing the targets. We thus decouple the distillation task from the MAE reconstruction task by connecting the decoder for reconstruction to other intermediate layers of the student ViT-Tiny instead of the last layer during pre-training in Fig. 9. We conduct a sweep over the intermediate layer index ranging from 2 to 12 for the optimal decoupled distillation. The fine-tuning results on some selected datasets (see Tab. 5) show that connecting to the 6th, 8th, or 10th layer can all further improve the performance over the model based on D-MAE (i.e., the column with the encoder layer index of 12 for reconstruction in Tab. 5). We use the 8th encoder layer for MAE reconstruction in the subsequent experimental analysis and validation by default, and denote this pre-training with decoupled distillation as D2-MAE. It can be seen in Fig. 10 that the attention weights on the higher layers after D2-MAE pre-training exhibit very similar behavior patterns to the ViT-Base teacher (see Fig. 7).
In Tab. 5, we also test the scenario of only relying on the attention map distillation to achieve good representation for the ViT-Tiny model, i.e., the MAE reconstruction loss in Fig. 9 is abandoned. It is shown that the experimental results without reconstruction are significantly lower than using both the distillation and reconstruction. However, only relying on the attention map distillation has already surpassed the MAE pre-training without distillation. This further demonstrates the effectiveness and superiority of our approach for developing the attention map distillation-based solution.
Table 6: Distillation significantly improves performance on various tasks. Top-1 accuracy is reported for classification tasks, mIoU is for ADE20K segmentation, AUC is for LaSOT tracking, and AP is for COCO detection and segmentation (Det./Seg.).
@@
Flowers Pets Aircraft Cars CIFAR100 iNat18 IN1K ADE20K LaSOT COCO
@@
$\Delta$ to Supervised -0.4 -2.7 +6.5 +2.6 -0.6 +1.5 +2.2 +1.8 +2.0 +2.1/+2.0